# Configurable Multi-Port Memory Architecture for High-Speed Data Communication

### v1

@@ Abstract
Memory management is necessary with the increasing number of multi-connected AI devices and data bandwidth issues. For this purpose, high-speed multi-port memory is used. The traditional multi-port memory solutions are hard-bounded to a fixed number of ports for read or write operations. In this work, we proposed a pseudo-quad-port memory architecture. Here, ports can be configured (1-port, 2-port, 3-port, 4-port) for all possible combinations of read/write operations for the 6T static random access memory (SRAM) memory array, which improves the speed and reduces the bandwidth for data transfer. The proposed architecture improves the bandwidth of data transfer by 4$\times$. The proposed solution provides 1.3$\times$ and 2$\times$ area efficiency as compared to dual-port 8T and quad-port 12T SRAM. All the design and performance analyses are done using 65nm CMOS technology.
@@ I Introduction
Memory plays a crucial role in electronic systems. With the advancement of edge AI devices, efficient memory management is essential. Due to their complex computational needs, graphic, audio, and video processing require heterogeneous and multi-core processors. Multi-port memories have become vital for CPUs since multi-core processors demand significant data transfer capabilities. For edge AI applications such as smartwatches and autonomous cars, which receive or transfer data from various sources, reconfigurable multi-port memories are necessary for fast and efficient design. However, multi-port memories come with increased area and leakage costs. Traditionally, dual-port SRAMs use an 8T SRAM bitcell, which has a high area and leakage and faces contention issues when read word line (RWL) and write word line (WWL) is selected for the same bitcell. The write driver can disturb stored data by driving the write bitline. Similar problems can occur in most conventional dual-port memories, such as 8T, 10T, and 12T configurations. Additionally, this memory architecture can only function as a two-port memory, with one port fixed for reading and the other for writing. [1, 2, 3].
@@
  • The architecture provides an area-efficient solution to design multi-port memory architecture compared to other available work.
@@ II-A5 Clock Generator
Figure 3: Block diagram of clock generator. Here, simulation is shown for 4 cases
@@ II-B Functionality of the Architecture
Figure 6: Layout of the overall architecture of the proposed quad-port memory architecture.
@@ III Performance Analysis
The proposed architecture is designed with a 65nm CMOS technology node using Cadence Virtuoso and evaluated using post-layout simulations. Fig. 6 shows the layout of the overall architecture shown in Fig. 1. Here, the wrapper circuit increases the area overhead by only 8% for an SRAM macro of 16Kb. This is minimal compared to the area overhead due to the bitcell modification for 8T/12T/20T, etc., to add the additional ports. Fig. 5 shows the simulation waveform for a quad-port-enabled system. Simulation waveforms show all the internal signals of the clock generator. Here, for each clock of $CLK$, the architecture is scheduled to work as 4-port, 3-port, 2-port, and 1-port, respectively. Here, it can be observed that in a single clock period of external clock $CLK$, the $BACK$ signal generates the $N$, and $CLK2$ generates the $N-1$ clock pulses, which is then fed to the FSM to change the state of the port. Which changed the states of the ports through the multiplexer, and all four ports write/read the data to/from the SRAM macro. Also, at each transition of the $CLKP$, the states were initialized to the highest priority port. Table II compares the proposed work with other published work. Here, the main advantage of the proposed design is that it uses an efficient 6T SRAM cell, while other work uses a higher number of transistor-based memory cells. This adds to the huge area overhead, while the area overhead of the proposed wrapper is very minimal. Another advantage of the proposed work is that all ports can be configured for both read and write operations, while in other works, the ports were fixed for either read or write operations. The overall architecture works at the clock frequency of 250MHZ, while the memory access frequency becomes 4$\times$, i.e., 1GHz at 1.2V supply.
@@
Area ($mm^{2}$) 0.05 0.08 0.085

### v2

@@ Abstract
Memory management is necessary with the increasing number of multi-connected AI devices and data bandwidth issues. For this purpose, high-speed multi-port memory is used. The traditional multi-port memory solutions are hard-bounded to a fixed number of ports for read or write operations. In this work, we proposed a pseudo-quad-port memory architecture. Here, ports can be configured (1-port, 2-port, 3-port, 4-port) for all possible combinations of read/write operations for the 6T static random access memory (SRAM) memory array, which improves the speed and reduces the bandwidth for data transfer. The proposed architecture improves the bandwidth of data transfer by 4$\times$. The architecture provides 1.3$\times$ and 2$\times$ area efficiency compared to dual-port 8T and quad-port 12T SRAM.
@@ I Introduction
Memory plays a crucial role in electronic systems. With the advancement of edge AI devices, efficient memory management is essential. Due to their complex computational needs, graphic, audio, and video processing require heterogeneous and multi-core processors. Multi-port memories have become vital for CPUs since multi-core processors demand significant data transfer capabilities. For edge AI applications such as smartwatches and autonomous cars, which receive or transfer data from various sources, reconfigurable multi-port memories are necessary for fast and efficient design. Distributed arithmetic-based architectures also require multi-port SRAM circuits for high throughput design [1].
However, multi-port memories come with increased area and leakage costs. Traditionally, dual-port SRAMs use an 8T SRAM bitcell, which has a high area and leakage and faces contention issues when read word line (RWL) and write word line (WWL) is selected for the same bitcell. The write driver can disturb stored data by driving the write bitline. Similar problems can occur in most conventional dual-port memories, such as 8T, 10T, and 12T configurations. Additionally, this memory architecture can only function as a two-port memory, with one port fixed for reading and the other for writing. [2, 3, 4].
@@
  • The architecture provides an area-efficient solution to design multi-port memory architecture.
@@ II-A5 Clock Generator
Figure 3: Block diagram of clock generator.
@@ III Performance Analysis
The proposed architecture is designed with a 65nm CMOS technology node using Cadence Virtuoso and evaluated using post-layout simulations. Fig. 5 shows the layout of the overall architecture shown in Fig. 1. Here, the wrapper circuit increases the area overhead by only 8% for an SRAM macro of 16Kb. This is minimal compared to the area overhead due to the bitcell modification for 8T/12T/20T, etc., to add the additional ports. Fig. 6 shows the simulation waveform for a quad-port-enabled system. Simulation waveforms show all the internal signals of the clock generator. Here, for each clock of $CLK$, the architecture is scheduled to work as 4-port, 3-port, 2-port, and 1-port, respectively. Here, it can be observed that in a single clock period of external clock $CLK$, the $BACK$ signal generates the $N$, and $CLK2$ generates the $N-1$ clock pulses, which is then fed to the FSM to change the state of the port. Further, it changes the states of the ports through the multiplexer, and all four ports write/read the data to/from the SRAM macro. Also, at each transition of the $CLKP$, the states were initialized to the highest priority port. Table II compares the proposed work with other published work. Here, the main advantage of the proposed design is that it uses an efficient 6T SRAM cell, while other work uses a higher number of transistor-based memory cells. This adds to the huge area overhead, while the area overhead of the proposed wrapper is very minimal. The advantage of area efficiency is compared by memory density for fare comparison among different technology nodes. Table II also shows the improved power efficiency of the proposed architectures. Another advantage of the proposed work is that all ports can be configured for both read and write operations, while in other works, the ports were fixed for either read or write operations. The overall architecture works at the clock frequency of 250MHZ, while the memory access frequency becomes 4$\times$, i.e., 1GHz at 1.2V supply.
@@
Area ($mm^{2}$) 0.057 0.05 0.08 0.085
Memory Density ($Kb/mm^{2}$) 80 50 189
Power Consumption ($mW$) 5.79 1.2 1.8